OpenStudy (anonymous):

For the investment with interest compounded annually, find the final balance and the total interest earned. $500 for 2 years at 8% A. final balance: $540 interest earned: $40 B. final balance: $580 interest earned: $80 C. final balance: $583.20 interest earned: $83.20 D. final balance: $620 interest earned: $120

OpenStudy (koikkara):

Hmm.. Substituting this formula, What u get ? A(t) = P(1+(r/n))^(nt) A(2) = 500(1+(0.08/1))^(1*2) Well, How much u got , A(2)= ??

OpenStudy (anonymous):

?

OpenStudy (koikkara):

@eddiessocool hope the above Equation Helps.. How much you got , Calculate A(2)=??

OpenStudy (koikkara):

A(2) = 500(1+(0.08/1))^(1*2) ... A(2) = How much ??

OpenStudy (anonymous):

A(2) = 500(1.08)^2 WHATS THE ^

OpenStudy (koikkara):

well.. accurately, it comes as A(t) = P(1+(r/n))^(nt) A(2) = 500(1+(0.08/1))^(1*2) A(2) = 500(1.08)^2, which is Equal to A(2) = $583.20 ---( Solution)---- Total balance = $583.20 Total interest = $83.20 Wondering If you are satisfied with this answer, Please close this Question. Thank You ! Keep in touch with Open Study.. Bye !